# An Introduction to Single-Antenna Radio Astronomical Polarimetry

### v2

@@ C.5 Congruence Eigenmatrices
Given ${\bf A}=a\,\mbox{$\sigma$}_{0}+\underline{\mbox{$a$}}\cdot\underline{\mbox{$\sigma$}}\quad\text{and}\quad{\bf B}=b\,\mbox{$\sigma$}_{0}+\underline{\mbox{$b$}}\cdot\underline{\mbox{$\sigma$}},$ equation (C7) yields ${\bf A}{\bf B}=(ab+\underline{\mbox{$a$}}\cdot\underline{\mbox{$b$}})\mbox{$\sigma$}_{0}+(a\underline{\mbox{$b$}}+b\underline{\mbox{$a$}}+i\underline{\mbox{$a$}}\times\underline{\mbox{$b$}})\cdot\underline{\mbox{$\sigma$}}$ Therefore, ${\bf A}{\bf B}={\bf B}{\bf A}$ if and only if $\underline{\mbox{$a$}}\times\underline{\mbox{$b$}}=0$.
@@ C.6 Pure Mueller Matrices
$\begin{split}N_{\mu}^{\nu}&=\frac{1}{2}{\mbox{$\sigma$}_{\mu}}\,\mbox{$:$}\,{{\left({\bf J}\otimes{\bf J}^{\dagger}\right)}\,\mbox{$:$}\,{\mbox{$\sigma$}_{\nu}}}\\ &=\frac{1}{2}{\mbox{$\sigma$}_{\mu}}\,\mbox{$:$}\,{{\bf J}}\left({\mbox{$\sigma$}_{\nu}}\,\mbox{$:$}\,{{\bf J}^{\dagger}}\right)\\ &=\frac{1}{2}k_{\mu}k^{*}_{\nu}.\end{split}$ (C34)
Here, $k_{\mu}={\mbox{$\sigma$}_{\mu}}\,\mbox{$:$}\,{{\bf J}}$ are the components of a complex-valued target four-vector $\underline{\underline{\mbox{$k$}}}$ (Cloude, 1986). Equation (C34) is equivalent to $\underline{\underline{{\bf N}}}=\underline{\underline{\mbox{$k$}}}\otimes\underline{\underline{\mbox{$k$}}}^{\dagger}/2$ and, as noted by Simon (1982), N is a scalar multiple of a projection matrix, such that
@@
The mapping between any Mueller matrix $\underline{\underline{{\bf M}}}$ and its associated target coherency matrix $\underline{\underline{{\bf N}}}$ exploits the symmetry of the transpose used to define the $\tilde{\otimes}$ operator (eqn. 45). Let $\mathcal{T}$ represent the transpose of covariant tensor indeces, such that ${\bf A}\,\tilde{\otimes}\,{\bf B}=\mathcal{T}({\bf A}\otimes{\bf B})$ and, by symmetry, ${\bf A}\otimes{\bf B}=\mathcal{T}({\bf A}\,\tilde{\otimes}\,{\bf B})$. Using this operator, the rank 4 tensors associated with M and N are related by $\mbox{$\mathcal{U}$}=\mathcal{T}(\mbox{$\mathcal{N}$})$ and
@@
Equation (C37) can be written as $\underline{\underline{{\bf N}}}=\underline{\underline{{\bf\mbox{$\mathcal{T}$}}}}:\underline{\underline{{\bf M}}}$, where $\underline{\underline{{\bf\mbox{$\mathcal{T}$}}}}$ is the four-dimensional rank 4 tensor defined by
@@
By symmetry, $\underline{\underline{{\bf M}}}=\underline{\underline{{\bf\mbox{$\mathcal{T}$}}}}:\underline{\underline{{\bf N}}}$; that is, $\underline{\underline{{\bf\mbox{$\mathcal{T}$}}}}$ is an involution. As for the target coherency matrix derived from a Jones matrix, $\underline{\underline{{\bf N}}}$ is Hermitian.
@@
Just as the coherency matrix formed by the outer product of a single Jones vector $\tilde{\mbox{$\rho$}}=\hat{\mbox{$e$}}\otimes\hat{\mbox{$e$}}^{\dagger}$ corresponds to a purely polarized state (see eqn. C17), the target coherency matrix formed by the outer product of a single target vector $\underline{\underline{{\bf N}}}=\underline{\underline{\mbox{$k$}}}\otimes\underline{\underline{\mbox{$k$}}}^{\dagger}/2$ corresponds to a pure Mueller matrix. Therefore, a Mueller matrix is pure if and only if its target coherency matrix, $\underline{\underline{{\bf N}}}=\underline{\underline{{\bf\mbox{$\mathcal{T}$}}}}:\underline{\underline{{\bf M}}}$ satisfies equation (C35). Cloude (1986) expressed this definition with the equivalent constraint that a pure Mueller matrix is associated with a target coherency matrix that has only one non-zero eigenvalue, $\lambda$. Combined with its associated eigenvector, $\hat{\underline{\underline{\mbox{$k$}}}}$, the target vector $\underline{\underline{\mbox{$k$}}}=\lambda^{1/2}\hat{\underline{\underline{\mbox{$k$}}}}$ can be used to compute (up to arbitrary phase) the Jones matrix associated with a pure Mueller matrix,
@@ C.7 Impure Mueller Matrices
where the angle brackets represent integration over time and $d=\exp(-2\sigma_{\Psi}^{2})<1$ describes the depolarization of Stokes Q and U. Note that this Mueller matrix is able to depolarize a purely polarized state, something that cannot be done by a linear transformation of the electric field (Proof 4.3).
@@ Appendix D Example Transformations
$\underline{\underline{\mbox{$S$}}}=[S_{0},S_{1},0,0]$ (D1)
@@ D.1 Example Boost
Consider a boost along $\hat{\underline{\mbox{$1$}}}=(1,0,0)$,
@@
$\begin{split}S_{0}^{\prime}&=\mbox{$\sigma$}_{0}:\mbox{$\rho$}^{\prime}=S_{0}\cosh 2\beta+S_{1}\sinh 2\beta\\ S_{1}^{\prime}&=\mbox{$\sigma$}_{1}:\mbox{$\rho$}^{\prime}=S_{0}\sinh 2\beta+S_{1}\cosh 2\beta.\end{split}$ (D7)
@@ D.2 Example Rotation
Consider a rotation about $\hat{\underline{\mbox{$2$}}}=(0,1,0)$,
@@
$\begin{split}S_{0}^{\prime}&=\mbox{$\sigma$}_{0}:\mbox{$\rho$}^{\prime}=S_{0}\\ S_{1}^{\prime}&=\mbox{$\sigma$}_{1}:\mbox{$\rho$}^{\prime}=S_{1}\cos 2\chi\\ S_{3}^{\prime}&=\mbox{$\sigma$}_{3}:\mbox{$\rho$}^{\prime}=S_{1}\sin 2\chi.\end{split}$ (D13)
@@ D.3 Complex-valued Polarization Ellipse
As in Section 5.3.2, consider the transformation from the original basis to one in which the second component of the electric field vector is zero; i.e.,
@@ D.4 Transformation to a Circular Basis
Equation (84) is not the only way to represent a pair of orthonormal circularly-polarized receptors; however, it is the only unimodular unitary matrix that effects the desired cyclic permutation of $\underline{\mbox{$S$}}=(Q,U,V)$ into $\underline{\mbox{$S$}}^{\prime}=(V,Q,U)$ for all values of $Q$, $U$ and $V$ (assuming the Stokes V sign convention of IEEE, 1983). To see this, consider the following three derivations of equation (84), each of which provides a slightly different perspective.
@@ D.4.1 Direct Derivation
As discussed in Section 3.4, for a left-hand circularly polarized (LCP) wave, the phase of $e_{y}$ leads that of $e_{x}$ by $90^{\circ}$. If the LCP wave is also 100% polarized, then
@@
$\hat{\mbox{$r$}}_{L}=\frac{1}{\sqrt{2}}\left(1\;\;\;-i\right)\quad\text{and}\quad\hat{\mbox{$r$}}_{R}=\frac{1}{\sqrt{2}}\left(1\;\;\;\;i\right).$ (D37)
@@ D.4.2 Indirect Derivation
$\hat{\mbox{$r$}}^{\prime}_{L}=\hat{\mbox{$r$}}_{L}=\frac{1}{\sqrt{2}}\left(1\;\;\;-i\right)$ (D39)
@@
$\hat{\mbox{$r$}}^{\prime}_{R}=-i\hat{\mbox{$r$}}_{R}=\frac{1}{\sqrt{2}}\left(-i\;\;\;\;1\right).$ (D40)
@@
rotates the polarization vector $\underline{\mbox{$S$}}=(Q,U,V)$ around the $S_{2}$ axis by $90^{\circ}$, yielding $\underline{\mbox{$S$}}^{\prime}=(V,U,-Q)$.
To arrive at the desired cyclic permutation, note that multiplying the receptors by different phases results in a rotation about the $S_{1}$ axis (see Section 5.2.2), and that a $90^{\circ}$ rotation about this axis transforms $\underline{\mbox{$S$}}^{\prime}=(V,U,-Q)$ into $\underline{\mbox{$S$}}^{\prime\prime}=(V,Q,U).$ Therefore,
@@ D.4.3 Geometric Derivation
Rather than starting with the electric field, consider the three-dimensional rotation of the Stokes polarization vector that effects the desired cyclic permutation. The eigenvectors of this rotation lie on the axis defined by $Q=U=V$; therefore, choose $\hat{\underline{\mbox{$n$}}}=(1,1,1)/\sqrt{3}$ and note that a $120^{\circ}$ rotation about this axis rotates Stokes $V$ into $S^{\prime}_{1}$, Stokes $Q$ into $S^{\prime}_{2}$, and Stokes $U$ into $S^{\prime}_{3}$. The equivalent unitary transformation of the electric field,
@@ Appendix E Stokes Parameters as Intensity Differences
Currently, at frequencies above the microwave region of the electromagnetic spectrum, it is not possible to directly sample the electric field and therefore not possible to simultaneously compute all four Stokes parameters as in equations (17) through (20). Rather, only $S_{0}$ and $S_{1}$ can be directly measured because they are the sums and differences of the flux densities (or intensities) of the radiation in two orthogonally polarized receptors.
In the basis defined by the Cartesian coordinates introduced in Section 2, the Stokes polarization vector $\underline{\mbox{$S$}}_{q}=(Q,U,V)$ and equation (18) yields
@@
In a basis that is formed by rotating the original basis by $45^{\circ}$ around the line of sight (e.g., the $x^{\prime}$ and $y^{\prime}$ axes of Section 2 when $\psi=\pi/4$), the Stokes polarization vector $\underline{\mbox{$S$}}_{u}=(U,-Q,V)$ and equation (18) yields
@@
In a basis defined by receptors with orthogonal senses of circular polarization, as discussed in Section 5.4.1, the Stokes polarization vector $\underline{\mbox{$S$}}_{v}=(V,Q,U)$, and equation (18) yields
@@ Appendix F Refinements to Existing Definitions
This article introduces some minor refinements and corrections to existing definitions that enhance their clarity and accuracy. First, in Stokes (1852), the ellipticity angle $\chi$ varies between $90^{\circ}$ and $90^{\circ}$, as inferred from “the numerical value of $\chi$] being supposed not to lie beyond the limits 0 and 90” and “polarization is right-handed or left-handed according to the sign of $\chi$].” (The “numerical value” is understood to mean the “absolute value.”) However, as noted in Section 7.3, allowing $|\chi|>45^{\circ}$ leads to model degeneracy; therefore $|\chi|\leq 45^{\circ}$ in this work.
@@ Appendix G Down-Conversion
The spectral information is shifted to baseband by demodulating or mixing the radio frequencies (RF) with a local oscillator (LO). This is equivalent to multiplying the signal, $x(t)$, with a pure tone, $l(t)=a\cos(2\pi\nu_{l}t+\phi)$. By application of the convolution theorem, and reference to Table 2, mixing may also be understood as a convolution with a pair of complex-valued delta functions in the frequency domain. This understanding proves useful in the following sections.
@@ G.1 Single-Sideband Down-Conversion
Figure 6: Upper-sideband down-conversion. From the top, the real-valued signal, $x(t)$ has a complex-valued spectrum with conjugate symmetry, $X(-\nu)=X^{*}(\nu)$, as depicted by $X(\nu)]$ in blue and ${\rm Im}[X(\nu)]$ in red. The signal is bandpass filtered and the band-limited signal, $X_{b}(\nu)$, is mixed with a local oscillator with frequency $\nu_{0}-\Delta\nu/2$, as depicted by a pair of blue delta functions. The resulting signal, $X_{m}(\nu)$, is low-pass filtered, producing $X_{u}(\nu)$. For each spectrum, the frequency axis is drawn in black, and the real and imaginary parts of each complex value are projected along the blue and red axes (respectively) depicted at the origin (where $\nu=0$).

### v4

@@ C.5 Congruence Eigenmatrices
Given ${\bf A}=a\,\mbox{$\sigma$}_{0}+\underline{\mbox{$a$}}\cdot\underline{\mbox{$\sigma$}}\quad\text{and}\quad{\bf B}=b\,\mbox{$\sigma$}_{0}+\underline{\mbox{$b$}}\cdot\underline{\mbox{$\sigma$}},$ Equation (C7) shows that ${\bf A}{\bf B}={\bf B}{\bf A}$ if and only if $\underline{\mbox{$a$}}\times\underline{\mbox{$b$}}=0$.
@@ C.6 Pure Mueller Matrices
$\begin{split}N_{\mu}^{\nu}&=\frac{1}{2}{\mbox{$\sigma$}_{\mu}}\mathbin{:}{{\left({\bf J}\otimes{\bf J}^{\dagger}\right)}\mathbin{:}{\mbox{$\sigma$}_{\nu}}}\\ &=\frac{1}{2}{\mbox{$\sigma$}_{\mu}}\mathbin{:}{{\bf J}}\left({\mbox{$\sigma$}_{\nu}}\mathbin{:}{{\bf J}^{\dagger}}\right)\\ &=\frac{1}{2}k_{\mu}k^{*}_{\nu}.\end{split}$ (C34)
Here, $k_{\mu}={\mbox{$\sigma$}_{\mu}}\mathbin{:}{{\bf J}}$ are the components of a complex-valued target four-vector $\underline{\underline{\mbox{$k$}}}$ (Cloude, 1986). Equation (C34) is equivalent to $\underline{\underline{{\bf N}}}=\underline{\underline{\mbox{$k$}}}\otimes\underline{\underline{\mbox{$k$}}}^{\dagger}/2$ and, as noted by Simon (1982), N is a scalar multiple of a projection matrix, such that
@@
The mapping between any Mueller matrix $\underline{\underline{{\bf M}}}$ and its associated target coherency matrix $\underline{\underline{{\bf N}}}$ exploits the symmetry of the transpose used to define the $\tilde{\otimes}$ operator (Eqn. 45). Let $\mathcal{T}$ represent the transpose of covariant tensor indeces, such that ${\bf A}\,\tilde{\otimes}\,{\bf B}=\mathcal{T}({\bf A}\otimes{\bf B})$ and, by symmetry, ${\bf A}\otimes{\bf B}=\mathcal{T}({\bf A}\,\tilde{\otimes}\,{\bf B})$. Using this operator, the rank 4 tensors associated with M and N are related by $\mbox{$\mathcal{U}$}=\mathcal{T}(\mbox{$\mathcal{N}$})$, where $\mathcal{U}$ is defined by Equation (50) and
@@
Equation (C37) can be written as $\underline{\underline{{\bf N}}}=\underline{\underline{{\bf\mbox{$\mathcal{T}$}}}}\mathbin{:}\underline{\underline{{\bf M}}}$, where $\underline{\underline{{\bf\mbox{$\mathcal{T}$}}}}$ is the four-dimensional rank 4 tensor defined by
@@
By symmetry, $\underline{\underline{{\bf M}}}=\underline{\underline{{\bf\mbox{$\mathcal{T}$}}}}\mathbin{:}\underline{\underline{{\bf N}}}$; that is, $\underline{\underline{{\bf\mbox{$\mathcal{T}$}}}}$ is an involution. As for the target coherency matrix derived from a Jones matrix, $\underline{\underline{{\bf N}}}$ is Hermitian.
@@
Just as the coherency matrix formed by the outer product of a single Jones vector $\tilde{\mbox{$\rho$}}=\hat{\mbox{$e$}}\otimes\hat{\mbox{$e$}}^{\dagger}$ corresponds to a purely polarized state (see Eqn. C17), the target coherency matrix formed by the outer product of a single target vector $\underline{\underline{{\bf N}}}=\underline{\underline{\mbox{$k$}}}\otimes\underline{\underline{\mbox{$k$}}}^{\dagger}/2$ corresponds to a pure Mueller matrix. Therefore, a Mueller matrix is pure if and only if its target coherency matrix, $\underline{\underline{{\bf N}}}=\underline{\underline{{\bf\mbox{$\mathcal{T}$}}}}\mathbin{:}\underline{\underline{{\bf M}}}$ satisfies Equation (C35). Cloude (1986) expressed this definition with the equivalent constraint that a pure Mueller matrix is associated with a target coherency matrix that has only one non-zero eigenvalue, $\lambda$. Combined with its associated eigenvector, $\hat{\underline{\underline{\mbox{$k$}}}}$, the target vector $\underline{\underline{\mbox{$k$}}}=\lambda^{1/2}\hat{\underline{\underline{\mbox{$k$}}}}$ can be used to compute (up to arbitrary phase) the Jones matrix associated with a pure Mueller matrix,
@@ C.7 Impure Mueller Matrices
where the angle brackets represent integration over time and $d=\exp(-2\sigma_{\Psi}^{2})<1$ describes the depolarization of Stokes Q and U. Note that this Mueller matrix is able to depolarize a purely polarized state, something that cannot be done by a linear transformation of the electric field (Proof IV.3).
@@ Appendix D Example Transformations
$\underline{\underline{\mbox{$S$}}}=[S_{0},S_{1},0,0]^{T}$ (D1)
@@ D.1 Example Boost
Consider a boost along $\hat{\underline{\mbox{$1$}}}=(1,0,0)^{T}$,
@@
$\begin{split}S_{0}^{\prime}&=\mbox{$\sigma$}_{0}\mathbin{:}\mbox{$\rho$}^{\prime}=S_{0}\cosh 2\beta+S_{1}\sinh 2\beta\\ S_{1}^{\prime}&=\mbox{$\sigma$}_{1}\mathbin{:}\mbox{$\rho$}^{\prime}=S_{0}\sinh 2\beta+S_{1}\cosh 2\beta.\end{split}$ (D7)
@@ D.2 Example Rotation
Consider a rotation about $\hat{\underline{\mbox{$2$}}}=(0,1,0)^{T}$,
@@
$\begin{split}S_{0}^{\prime}&=\mbox{$\sigma$}_{0}\mathbin{:}\mbox{$\rho$}^{\prime}=S_{0}\\ S_{1}^{\prime}&=\mbox{$\sigma$}_{1}\mathbin{:}\mbox{$\rho$}^{\prime}=S_{1}\cos 2\chi\\ S_{3}^{\prime}&=\mbox{$\sigma$}_{3}\mathbin{:}\mbox{$\rho$}^{\prime}=S_{1}\sin 2\chi.\end{split}$ (D13)
@@ D.3 Complex-valued Polarization Ellipse
As in Section V.3.2, consider the transformation from the original basis to one in which the second component of the electric field vector is zero; i.e.,
@@ D.4 Transformation to a Circular Basis
Equation (84) is not the only way to represent a pair of orthonormal circularly-polarized receptors; however, it is the only unimodular unitary matrix that effects the desired cyclic permutation of $\underline{\mbox{$S$}}=(Q,U,V)^{T}$ into $\underline{\mbox{$S$}}^{\prime}=(V,Q,U)^{T}$ for all values of $Q$, $U$ and $V$ (assuming the Stokes V sign convention of IEEE, 1983). To see this, consider the following three derivations of Equation (84), each of which provides a slightly different perspective.
@@ D.4.1 Direct Derivation
As discussed in Section III.4, for a left-hand circularly polarized (LCP) wave, the phase of $e_{y}$ leads that of $e_{x}$ by $90^{\circ}$. If the LCP wave is also 100% polarized, then
@@
$\hat{\mbox{$r$}}_{L}=\frac{1}{\sqrt{2}}\left(1,\,-i\right)\quad\text{and}\quad\hat{\mbox{$r$}}_{R}=\frac{1}{\sqrt{2}}\left(1,\,i\right).$ (D37)
@@ D.4.2 Indirect Derivation
$\hat{\mbox{$r$}}^{\prime}_{L}=\hat{\mbox{$r$}}_{L}=\frac{1}{\sqrt{2}}\left(1,\,-i\right)$ (D39)
@@
$\hat{\mbox{$r$}}^{\prime}_{R}=-i\hat{\mbox{$r$}}_{R}=\frac{1}{\sqrt{2}}\left(-i,\,1\right).$ (D40)
@@
rotates the polarization vector $\underline{\mbox{$S$}}=(Q,U,V)^{T}$ around the $S_{2}$ axis by $90^{\circ}$, yielding $\underline{\mbox{$S$}}^{\prime}=(V,U,-Q)^{T}$.
To arrive at the desired cyclic permutation, note that a $90^{\circ}$ rotation about the $S_{1}$ axis transforms $(V,U,-Q)^{T}$ into $(V,Q,U)^{T}.$ Therefore,
@@ D.4.3 Geometric Derivation
Rather than starting with the electric field, consider the three-dimensional rotation of the Stokes polarization vector that effects the desired cyclic permutation. The eigenvectors of this rotation lie on the axis defined by $Q=U=V$; therefore, choose $\hat{\underline{\mbox{$n$}}}=(1,1,1)^{T}/\sqrt{3}$ and note that a $120^{\circ}$ rotation about this axis rotates Stokes $V$ into $S^{\prime}_{1}$, Stokes $Q$ into $S^{\prime}_{2}$, and Stokes $U$ into $S^{\prime}_{3}$. The equivalent unitary transformation of the electric field,
@@ Appendix E Stokes Parameters as Intensity Differences
Currently, at frequencies above the microwave region of the electromagnetic spectrum, it is not possible to directly sample the electric field and therefore not possible to simultaneously compute all four Stokes parameters as in Equations (17) through (20). Rather, only $S_{0}$ and $S_{1}$ can be directly measured because they are the sums and differences of the flux densities (or intensities) of the radiation after passing through oppositely polarized filters.
In the basis defined by the Cartesian coordinates introduced in Section II, the Stokes polarization vector $\underline{\mbox{$S$}}_{q}=(Q,U,V)^{T}$, and Equation (18) yields
@@
In a basis that is formed by rotating the original basis by $45^{\circ}$ around the line of sight (e.g., the $x^{\prime}$ and $y^{\prime}$ axes of Section II when $\psi=\pi/4$), the Stokes polarization vector $\underline{\mbox{$S$}}_{u}=(U,-Q,V)^{T}$, and Equation (18) yields
@@
In a basis defined by receptors with orthogonal senses of circular polarization, as discussed in Section V.4.1, the Stokes polarization vector $\underline{\mbox{$S$}}_{v}=(V,Q,U)^{T}$, and Equation (18) yields
@@ Appendix F Refinements to Existing Definitions
This article introduces some minor refinements and corrections to existing definitions that enhance their clarity and accuracy. First, in Stokes (1852), the ellipticity angle $\chi$ varies between $90^{\circ}$ and $90^{\circ}$, as inferred from “the numerical value of $\chi$] being supposed not to lie beyond the limits 0 and 90” and “polarization is right-handed or left-handed according to the sign of $\chi$].” (The “numerical value” is understood to mean the “absolute value.”) However, as noted in Section VII.3, allowing $|\chi|>45^{\circ}$ leads to model degeneracy; therefore $|\chi|\leq 45^{\circ}$ in this work.
@@ Appendix G Down-conversion
The spectral information is shifted to baseband by demodulating or mixing the radio frequencies (RF) with a local oscillator (LO). This is equivalent to multiplying the signal, $x(t)$, with a pure tone, $l(t)=a\cos(2\pi\nu_{l}t+\phi)$. By application of the convolution theorem, and with reference to Table 2, mixing may also be understood as a convolution with a pair of complex-valued delta functions in the frequency domain. This understanding proves useful in the following sections.
@@ G.1 Single-sideband Down-conversion
Figure 6: Upper-sideband down-conversion. From the top, the real-valued signal, $x(t)$ has a complex-valued spectrum with conjugate symmetry, such that $X(-\nu)=X^{*}(\nu)$, as depicted by $X(\nu)]$ in blue and ${\rm Im}[X(\nu)]$ in red. The signal is bandpass filtered and the band-limited signal, $X_{b}(\nu)$, is mixed with a local oscillator with frequency $\nu_{0}-\Delta\nu/2$, as depicted by a pair of blue delta functions. The resulting signal, $X_{m}(\nu)$, is low-pass filtered, producing $X_{u}(\nu)$. For each spectrum, the frequency axis is drawn in black, and the real and imaginary parts of each complex value are projected along the blue and red axes (respectively) depicted at the origin (where $\nu=0$).